Under the action of load F1, the length of a string is L1 and that under F2, is L2. the original length of the wire is

(a) [L1F1 - L2F2] / [F1 + F2]

(b) [L1F2 - L2F1] / [F1 - F2]

(c) [L1F2 - L2F1] / [F2 - F1]

(d) [L1F2 - L2F1] / [F1 + F2]

Grade:11

2 Answers

Arun
3 years ago
Dear Jayant
 

Let dL be the change in length

and L be the Original length

Therefore Y=(f1*L)/(A*dL) for L1

= (f1*L)/A*(L-L1)--------(1)

Y=(f2*L)/(A*dL) forL2

=(f2*L)/A(l-L2)-------(2)

Equating both 1&2

F1/(L-L1)=f2/(L-L2)

=>L=(f1L2-f2L1)/(f1-f2)

 

Hope it helps

Vikas TU
3 years ago
Dear student 
Question is not clear 
Please attach an image, 
We will happy to  help you 
Good Luck
Cheers